There are 7 boys and 5 girls in a youngsters’ club. A committee of 3 boys and 2 girls is to be chosen. How many different possibilities are there?

1
Expert's answer
2022-03-22T13:57:31-0400

"N=C_7^3*C_5^2=35*10=350."